OpenStudy (anonymous):

SR18+ SR 4y^3-2SR y^3 -SR 48y^2 solve this radical!

OpenStudy (anonymous):

Try once

OpenStudy (anonymous):

question is \[\sqrt{18} + \sqrt{4y^3} -2\sqrt{y^3}-\sqrt{48y^2}\]

OpenStudy (anonymous):

ive been working them out, im checking my answer which i got. 2y^2+3y SR2y. is that right?

OpenStudy (anonymous):

no it is not correct

OpenStudy (anonymous):

because first term is \[\sqrt{18}= 3\sqrt{2}\]

OpenStudy (radar):

\[\sqrt{18}+\sqrt{4y ^{3}}-2\sqrt{y ^{3}}-\sqrt{48y ^{2}}\] Do them one at a time, slowly and carefully.

OpenStudy (radar):

sheg, has worked the first one, look at the next two, and figure out why you can just throw them away !

OpenStudy (anonymous):

@riyamoon it is still not correct If your helping me with my question its not letting me post. i got \[3\sqrt{2}+6\sqrt{3}\]

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

no it is not correct

OpenStudy (anonymous):

my computer is being gay...then idk what im doing wrong

OpenStudy (radar):

The last term can be also expressed as\[\sqrt{16 * 3 y ^{2}}\] Does that help?

OpenStudy (radar):

Note that two of those are perfect squares!

OpenStudy (radar):

Now you can work it

OpenStudy (anonymous):

@riyamoon are you trying to throwing the answer blindly without doing the calculations

OpenStudy (radar):

@riyamoon, what is the square root of 16, what is the square root of \[y ^{2}????\]

OpenStudy (anonymous):

riya u need to work hard

OpenStudy (radar):

Looking at each term of the expressionL\[\sqrt{18}=\sqrt{9}\sqrt{2}=3\sqrt{2}\] Looking at the next two,\[+\sqrt{4y ^{3}}-2\sqrt{y ^{3}}=+2\sqrt{y ^{3}}-2\sqrt{y ^{3}}=0\] That leaves the last one for you, and you were shown how it can be broken down and solved.

OpenStudy (anonymous):

\[\sqrt{18} +\sqrt{4y^3} - 2\sqrt{y^3} - \sqrt{48y^2}\] \[=3\sqrt{2} +2y\sqrt{y} - 2y\sqrt{y} - 4y\sqrt{3}\] \[=3\sqrt{2} - 4y\sqrt{3}\]